2. Citizens of the country of Heehaw produce hay and provide entertainment services (banjo playing). In one year they produced $15 million worth of hay, with $11 million consumed domestically and the other $4 million sold to neighboring countries. They provided $7 million worth of banjo-playing services, $5 million in Heehaw, and $2 million in neighboring countries. They purchased $6 million worth of soda pop from neighboring countries. Calculate the magnitudes of GNP, GDP, net factor payments from abroad, net exports, and the current account balance.
Answer #1

GNP= output by citizens=15+7=22 million$

GDP= output produced by country =15(Hay)+5(donestic banjo) =20million$

Net factor payment = GNP - GDp= 2million$

Net exports= 4(Hay sold abroad) - 6(soda imports) = - 2million$

Current account balance = netexports +Net factor payment = - 2+2= 0.